# Correlated phases of moat-band excitons in two-dimensional systems

### v1

@@ IX Acknowledgments
This work is supported by the Delta-ITP consortium, which is part of the Netherlands Organisation for Scientific Research (NWO). We also acknowledge the research program “Materials for the Quantum Age” (QuMat) for financial support. This program (registration number 024.005.006) is part of the Gravitation program financed by the Dutch Ministry of Education, Culture and Science (OCW).
@@ Appendix D Interaction energy of the chiral spin liquid
where $\ell_{n}\equiv(2\pi n)^{{-}1/2}$ is the magnetic length. Either expression results in Eq. (24) of the main text. for the bilayer interaction of Eq. (8). The first term gives the zero-point energy $(n/2)\int\mathrm{d}^{2}r\,V(r)$, while the second term corresponds to the actual correlation energy of the usual QHE. In general, if the density is exactly $\rho_{l}\equiv{-}\lambda/2(2l+1)$ so that the kinetic energy is perfectly quenched, the interaction energy in the case of the bilayer potential is given by
@@
where the $\rho^{2}_{l}$ term cancels because $\int_{0}^{\infty}\mathrm{d}x\,[L_{l}(x)]^{2}\hskip 0.85358pt\mathrm{e}^{-x}=1$. Eq. (74) shows that, in the case of a perfectly degenerate moat, the CSL will always have a lower energy than the BEC phases at small densities, as $\tilde{E}_{\text{int}}^{\mathrm{(CSL)}}\propto\rho^{3}$ while $\tilde{E}_{\text{int}}^{\mathrm{(FF)}}\propto\rho^{2}$. Furthermore, the rightmost LL is that with $l=0$, so there is a maximum value of $\rho$ for which the LL energy can be nullified and above which the BEC phases can become stable. In practice, as explained in Sec. III.4, band-structure warping in real systems will likely destabilize the CSL phase in a large region of the parameter space.

### v2

@@ IX Acknowledgments
This work is supported by the Delta-ITP consortium, which is part of the Netherlands Organisation for Scientific Research (NWO). We also acknowledge the research program “Materials for the Quantum Age” (QuMat) for financial support. This program (registration number 024.005.006) is part of the Gravitation program financed by the Dutch Ministry of Education, Culture and Science (OCW). We thank Tigran Sedrakyan and Rui Wang for valuable comments and useful discussions.
@@ Appendix D Interaction energy of the chiral spin liquid
where $\ell_{n}\equiv(2\pi n)^{{-}1/2}$ is the magnetic length. Either expression results in Eq. (24) of the main text. for the bilayer interaction of Eq. (8). The first term gives the zero-point energy $(n/2)\int\mathrm{d}^{2}r\,V(r)$, while the second term corresponds to the actual correlation energy of the usual QHE. In general, if the density is exactly $\rho_{l}\equiv{-}\lambda/[2(2l+1)]$ so that the kinetic energy is perfectly quenched, the interaction energy in the case of the bilayer potential is given by
@@
where the $\rho^{2}_{l}$ term cancels because $\int_{0}^{\infty}\mathrm{d}x\,[L_{l}(x)]^{2}\hskip 0.85358pt\mathrm{e}^{-x}=1$. Eq. (74) shows that, in the case of a perfectly degenerate moat, the CSL will always have a lower energy than the BEC phases at small densities, as $\tilde{E}_{\text{int}}^{\mathrm{(CSL)}}\propto\rho^{3}$ while $\tilde{E}_{\text{int}}^{\mathrm{(FF)}}\propto\rho^{2}$. Furthermore, the rightmost LL is that with $l=0$, so there is a maximum value of $\rho$ for which the LL energy can be nullified and above which the BEC phases can become stable. In practice, as explained in Sec. III.4, band-structure warping in certain systems can drive a transition from the CSL to the BEC already at lower densities.